OpenStudy (anonymous):

What is the distance between the points (-6, -2) and (2, 4)? a. 2 √5 b. 2 √7 c. 10 d. 28

OpenStudy (anonymous):

10 units

OpenStudy (anonymous):

c

OpenStudy (anonymous):

d=sqrt[(2+6)^2+(4+2)^2] d=sqrt(64+36) d=sqrt100 d=10 c option is correct

OpenStudy (anonymous):

\[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

distance formula